1. Given the following static system, find the required value of F. 8N 6N
Answer #1


x axis is to the right and y axis is upward.

so F1 = 8i and F2 = - 6j

for static system, Fnet = F1 + F2 + F = 0

F = - 8i + 6j

magnitude = sqrt(8^2 + 6^2) = 10 N

direction = tan^-1(6/8) = 37 deg


So 10 N at 37 deg above from the left side.